OpenStudy (anonymous):

3x2 − 3y + 8x − 5x2 + 12y

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

Simplify the expression

OpenStudy (anonymous):

are those x's or a multiplication symbol

OpenStudy (anonymous):

x's

OpenStudy (anonymous):

ok, just checking. You have (3*x*2)-3y +8x - (5*x*2) +12y for the equation then

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the first thing you can do is simplify whats in parenthesis. 3*x*2 = 6x 5*x*2 = 10x

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

then group together like terms starting with 6x - 3y +8x - 10y + 12y = (6x + 8x) + (-3y -10y +12y)

OpenStudy (anonymous):

from there just add the terms. 14x -1y

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

if sorry i realized thats supposed to be 10x not 10y

OpenStudy (anonymous):

i got it im a little slow in math. but thanks

OpenStudy (anonymous):

its really (6x+8x-10x)+(-3y+12y) = 4x+9y